OpenStudy (anonymous):

solve 5x + 12 = x -4

OpenStudy (anonymous):

5x-x=-12-4

OpenStudy (anonymous):

4x=-16 so x=-16/4=-4

OpenStudy (compassionate):

\(\Huge\color{blue}{\mathbb{Welcome~To~OpenStudy!!!}}\) 5x + 12 = x -4 Add x to the left side. 5x + x + 12 = -4 6x + 12 = -4 Subtract 12 6x = -16 Now divide \[\frac{-16}{6}\]

OpenStudy (compassionate):

Whoops, you want to subtact x!!! It should be: 4x = -16

OpenStudy (anonymous):



OpenStudy (anonymous):

yes @Compassionate i was about to point that out to you :) lol but its okaii like you always say, it happens to even the best of us ;)